OpenStudy (anonymous):

one leg of a right triangle is three times the length of the other leg. the hypotenuse is 20 inches long. find the length of each leg?

OpenStudy (yttrium):

20^2 = x^2+9x^2 400 = 10x^2 40 = x^2 x = 2sqrt10 -- length of the shorter leg 3x = 6sqrt10 --- length of the longer leg

OpenStudy (anonymous):

omg thanks so much

OpenStudy (whpalmer4):

a^2 + b^2 = c^2 in a right triangle, where c is the length of the hypotenuse, here 20. a^2 + b^2 = 20^2 = 400 we know that a = 3b because one leg is three times the length of the other. a^2 + b^2 = 400 (3b)^2 + b^2 = 400 9b^2 + b^2 = 400 10b^2 = 400 b^2 = 40 b = sqrt 40 = sqrt 4 * sqrt 10 = 2 sqrt 10 a = 3b = 3*2 sqrt 10 = 6 sqrt 10